# Mosaic: Composite Projection Pruning for Resource-efficient LLMs

### v1

@@ V-B3 Composite Projection Pruning Performance (E3)
V-B4 Fine-tuning Performance (E4)
The SLM quality is regained after pruning using parameter-efficient fine-tuning (PEFT) approaches, such as low-rank adaptation [28] (LoRA). Alpaca [41] (Table III) is used to train a LoRA adapter for LLaMa-3.1-8B for two epochs [13]. LoRA creates an 84 MB adapter which merges into the original pruned model weights at runtime. Figure 10 and Table VI summarize the findings.
@@ V-B4 Fine-tuning Performance (E4)
Full Dataset Evaluation: Table VI demonstrates that by fine-tuning on one dataset, namely Alpaca, the perplexity and accuracy improves on other datasets listed in Table III. All methods show better perplexity and accuracy following fine-tuning. Projection pruning starts with higher accuracy and gains more from fine-tuning than other methods. This indicates that projection pruning retains an optimal set of parameters that can be effectively retrained.
@@ V-B5 End-to-End Overheads (E5)
The pruning overhead (purple bars) of layer and projection pruning is higher than global pruning. This is because a weight metric is calculated for every parameter to obtain their pruning targets (Section III-A), then prune the LLM based on target percentages. Global pruning of LLaMa-3.1-8B and LLaMa-2-13B requires 23.31 GB and 33.34 GB of memory, respectively, for the model parameters in FP16 precision, software libraries, 128 samples from the calibration dataset, and model activations from calibration. Layer and projection pruning requires additional memory due to the weight metrics - 24.22 GB and 35.81 GB for LLaMa-3.1-8B and LLaMa-2-13B, respectively.
@@
As explored in Section V-B4, LLMs are typically fine-tuned for deployment to regain accuracy lost during pruning. The fine-tuning overhead is shown in Figure 11 (orange bars). The models pruned using layer and projection pruning are fine-tuned to match the same accuracy achieved by global pruning after it is fine-tuned for two epochs. For LLaMa-3.1-8B, projection pruning produces a deployment ready model over 4.8 faster than global and layer pruning. In the case of LLaMa-2-13B, projection pruning is 2.67 and 7.19 faster than layer and global pruning, respectively.
VI Related Work
@@
Compressing LLMs: The three main LLM compression methods to improve resource efficiency are model quantization, pruning, and knowledge distillation.
@@
Model Quantization - State-of-the-art quantization methods, such as GPTQ [29] and AWQ [44] reduce FP16 weights of the LLM into lower precision, such as INT3, INT4, or INT8, which requires less memory. Although the model size is reduced by more than half with minimal accuracy loss, the overall memory required remains high because model activations are not quantized and still use the higher FP16 precision. As memory demands grow due to larger activations from longer input lengths (more tokens), the quantized weights account for only a small fraction of the total memory usage, which reduces the overall effectiveness of this approach. Alternative methods that quantize model activations have lower accuracies and rely on a decomposition scheme to keep outlier weights at FP16 [45]. Although all quantization methods reduce memory requirements, orthogonal approaches are required to lower inference latency. For example, GPTQ relies on custom CUDA kernels [29], and AWQ utilizes QKV kernel fusion [44]. They are, therefore, suited to GPU-based systems for performance gains, which may not be available in resource-constrained environments.
Model Pruning - two categories of LLM pruning remove redundant parameters: unstructured pruning, which preserves model accuracy, and structured pruning, which reduces memory and computational requirements.
SparseGPT [12] and Wanda [21] are unstructured pruning methods. SparseGPT implements Optimal Brain Surgeon (OBS) [46] and prunes models with hundreds of billions of parameters in a few hours. In contrast, Wanda employs a simpler pruning metric based solely on model weights and activations, making it two orders of magnitude faster than SparseGPT. Unstructured pruning does not reduce model size since parameters are set to zero. These models require specialized sparse acceleration libraries, such as NVIDIA CUTLASS, which is limited to models that are 50% sparsity pruned using a specific semi-structured format [12, 21, 47].
LLM-Pruner [13] and Sheared-LLaMa [48] are structured pruning methods and produce smaller and faster LLMs. LLM-Pruner establishes pruning groups to efficiently prune LLM neurons in a one-shot manner, followed by fine-tuning to regain accuracy. In contrast, Sheared-LLaMa combines pruning and training in an iterative process. The pruned models are resource-efficient, but the overall model accuracy is lower than the original model. The above methods reduce layer size by removing neurons. BlockPruner [49] and ShortGPT [50] remove entire blocks and transformer layers, respectively.
@@
Knowledge Distillation - Smaller student models are trained to replicate the output of larger teacher models; they are used in older transformer architectures, such as BERT [2] with DistilBERT [51]. Minitron [14] is created using neural architecture search from Nemotron [52] using pruning and knowledge distillation. Although Minitron requires 40 fewer training tokens than other similar-sized LLMs to reach the same accuracy, the computational requirements are extensive (128 A100 80 GB GPUs [14]).
@@
Combining Unstructured and Structured Pruning for LLMs: Structured pruning of sparse models created by unstructured pruning has been explored for CNNs to make sparse models run on general-purpose hardware without the need for accelerators [26, 25]. They create small and fast sparse models. Mosaic is proposed to synergistically combine both pruning methods for LLMs and apply them at the granularity of projections rather than layers or blocks for the first time, referred to as composite projection pruning.
VII Conclusion
@@ Appendix B Model Sources
Table IX provides the Hugging Face model sources for use in the Transformers Python library for each LLM used in the experiments presented in the article. We provide the exact source as the same LLM can come in different formats, such as serialized Python objects (.bin) or Safetensors (.safetensors) depending on the Transformers library version used. In addition, the same LLM may come in FP16 or FP32 bit precision, depending on the source.
@@ Appendix C Detailed Accuracy Tables
Table X and Table XI provide the full zero-shot accuracy breakdown for LLaMa-3.1-8B and LLaMa-2-13B results for global, layer, and projection pruning. Note that at lower pruning targets, projection pruning does not outperform all other methods; however, on average, projection pruning achieves a higher accuracy. At higher pruning targets, projection pruning demonstrates superior performance.
@@ Appendix E Comparison to Older Pruning Methods
Table XII shows LLaMa-7B pruned by 70% zero-shot accuracy for all seven datasets compared to older LLM pruning methods. This model and pruning target are commonly used in the literature as most methods tend to collapse beyond the 70% pruning target, and LLaMa-7B was popular then. While more modern and better-trained LLMs are available now and were evaluated in the experiments, these tables provide a reference to compare older pruning methods.
@@ Appendix F Model Quantization
Model quantization is another compression method for LLMs that reduces the bit precision of the model parameters. This reduces the memory footprint of the model. Table XIII compares GPTQ, a quantization method used for LLMs, for four different bit targets using the group hyperparameter of 128 for LLaMa-3.1-8B compared against Mosaic. Zero-shot accuracy, as well as speedup and compression, is presented. Speedup is the improvement to LLM inference. While GPTQ can accelerate inference with custom kernels, the results presented are for hardware (P1) without the software accelerator. Mosaic does not require custom kernels to accelerate inference. Compression is the file size compression of the LLM weights. While GPTQ sets weights to the target bit size, activations are still 16 bits during inference, which has the same memory usage as the unquantized/unpruned (dense) model. For Mosaic, the pruning target proportionally decreases inference memory usage.

### v2

@@ 4.2.3 Composite Projection Pruning Performance (E3)
4.2.4 Fine-tuning Performance (E4)
The SLM quality is regained after pruning using parameter-efficient fine-tuning (PEFT) approaches, such as low-rank adaptation [28] (LoRA). Alpaca [41] (Table 3) is used to train a LoRA adapter for LLaMa-3.1-8B for two epochs [13]. LoRA creates an 84 MB adapter which merges into the original pruned model weights at runtime. Figure 10 and Table 6 summarize the findings.
@@ 4.2.4 Fine-tuning Performance (E4)
Full Dataset Evaluation: Table 6 demonstrates that by fine-tuning on one dataset, namely Alpaca, the perplexity and accuracy improves on other datasets listed in Table 3. All methods show better perplexity and accuracy following fine-tuning. Projection pruning starts with higher accuracy and gains more from fine-tuning than other methods. This indicates that projection pruning retains an optimal set of parameters that can be effectively retrained.
@@ 4.2.5 End-to-End Overheads (E5)
The pruning overhead (purple bars) of layer and projection pruning is higher than global pruning. This is because a weight metric is calculated for every parameter to obtain their pruning targets (Section 3.1), then prune the LLM based on target percentages. Global pruning of LLaMa-3.1-8B and LLaMa-2-13B requires 23.31 GB and 33.34 GB of memory, respectively, for the model parameters in FP16 precision, software libraries, 128 samples from the calibration dataset, and model activations from calibration. Layer and projection pruning requires additional memory due to the weight metrics - 24.22 GB and 35.81 GB for LLaMa-3.1-8B and LLaMa-2-13B, respectively.
@@
As explored in Section 4.2.4, LLMs are typically fine-tuned for deployment to regain accuracy lost during pruning. The fine-tuning overhead is shown in Figure 11 (orange bars). The models pruned using layer and projection pruning are fine-tuned to match the same accuracy achieved by global pruning after it is fine-tuned for two epochs. For LLaMa-3.1-8B, projection pruning produces a deployment ready model over 4.8 faster than global and layer pruning. In the case of LLaMa-2-13B, projection pruning is 2.67 and 7.19 faster than layer and global pruning, respectively.
5 Related Work
@@
The three main LLM compression methods to improve resource efficiency are model quantization, knowledge distillation, and pruning.
5.1 Model Quantization
@@
State-of-the-art quantization methods, such as GPTQ [29] and AWQ [44] reduce FP16 weights of the LLM into lower precision, such as INT3, INT4, or INT8, which requires less memory. Although the model size is reduced by more than half with minimal accuracy loss, the overall memory required remains high because model activations are not quantized and still use the higher FP16 precision. As memory demands grow due to larger activations from longer input lengths (more tokens), the quantized weights account for only a small fraction of the total memory usage, which reduces the overall effectiveness of this approach. Alternative methods that quantize model activations have lower accuracies and rely on a decomposition scheme to keep outlier weights at FP16 [45]. Although all quantization methods reduce memory requirements, orthogonal approaches are required to lower inference latency. For example, GPTQ relies on custom CUDA kernels [29], and AWQ utilizes QKV kernel fusion [44]. They are, therefore, suited to GPU-based systems for performance gains, which may not be available in resource-constrained environments.
5.2 Knowledge Distillation
@@
Smaller student models are trained to replicate the output of larger teacher models; they are used in older transformer architectures, such as BERT [2] with DistilBERT [46]. Minitron [14] is created using neural architecture search from Nemotron [47] using pruning and knowledge distillation. Although Minitron requires 40 fewer training tokens than other similar-sized LLMs to reach the same accuracy, the computational requirements are extensive (128 Nvidia A100 80 GB GPUs [14]).
5.3 Model Pruning
Unstructured Pruning: SparseGPT [12] and Wanda [21] are unstructured pruning methods. SparseGPT implements Optimal Brain Surgeon (OBS) [48] and prunes models with hundreds of billions of parameters in a few hours. In contrast, Wanda employs a simpler pruning metric based solely on model weights and activations, making it two orders of magnitude faster than SparseGPT. Unstructured pruning does not reduce model size since parameters are set to zero. These models require specialized sparse acceleration libraries, such as NVIDIA CUTLASS, which is limited to models that are 50% sparsity pruned using a specific semi-structured format [12, 21, 49].
Structured Pruning: LLM-Pruner [13] and Sheared-LLaMa [50] are structured pruning methods and produce smaller and faster LLMs. LLM-Pruner establishes pruning groups to efficiently prune LLM neurons in a one-shot manner, followed by fine-tuning to regain accuracy. In contrast, Sheared-LLaMa combines pruning and training in an iterative process. The pruned models are resource-efficient, but the overall model accuracy is lower than the original model. The above methods reduce layer size by removing neurons. BlockPruner [51] and ShortGPT [52] remove entire blocks and transformer layers, respectively.
@@ 5.3 Model Pruning
Combining Unstructured and Structured Pruning for LLMs: Structured pruning of sparse models created by unstructured pruning has been explored for CNNs to make sparse models run on general-purpose hardware without the need for accelerators [25, 26].
They create small and fast sparse models. Mosaic is proposed to synergistically combine both pruning methods for LLMs and apply them at the granularity of projections rather than layers or blocks for the first time, referred to as composite projection pruning. While existing work has explored composite pruning primarily in the context of CNNs [25, 26], Mosaic extends this approach to LLMs by targeting projection components unique to their architecture.
6 Conclusion
@@
This work has considered dense LLM architectures. Applying projection pruning to Mixture-of-Experts (MoE) models, such as Mixtral or DeepSeek, presents a promising direction for future research. MoE architectures introduce unique challenges, such as dynamic expert routing and load balancing, that may require new pruning strategies beyond those explored here. Extending projection pruning to MoEs remains an open area for continued investigation. Additionally, many LLM pruning methods rely on calibration datasets; future work will explore developing ranking methods that remove this dependency.
@@ Appendix B Model Sources
Table 9 provides the Hugging Face model sources for use in the Transformers Python library for each LLM used in the experiments presented in the article. We provide the exact source as the same LLM can come in different formats, such as serialized Python objects (.bin) or Safetensors (.safetensors) depending on the Transformers library version used. In addition, the same LLM may come in FP16 or FP32 bit precision, depending on the source.
@@ Appendix C Detailed Accuracy Tables
Table 10 and Table 11 provide the full zero-shot accuracy breakdown for LLaMa-3.1-8B and LLaMa-2-13B results for global, layer, and projection pruning. Note that at lower pruning targets, projection pruning does not outperform all other methods; however, on average, projection pruning achieves a higher accuracy. At higher pruning targets, projection pruning demonstrates superior performance.
@@ Appendix E Comparison to Older Pruning Methods
Table 12 shows LLaMa-7B pruned by 70% zero-shot accuracy for all seven datasets compared to older LLM pruning methods. This model and pruning target are commonly used in the literature as most methods tend to collapse beyond the 70% pruning target, and LLaMa-7B was popular then. While more modern and better-trained LLMs are available now and were evaluated in the experiments, these tables provide a reference to compare older pruning methods.
@@ Appendix F Model Quantization
Model quantization is another compression method for LLMs that reduces the bit precision of the model parameters. This reduces the memory footprint of the model. Table 13 compares GPTQ, a quantization method used for LLMs, for four different bit targets using the group hyperparameter of 128 for LLaMa-3.1-8B compared against Mosaic. Zero-shot accuracy, as well as speedup and compression, is presented. Speedup is the improvement to LLM inference. While GPTQ can accelerate inference with custom kernels, the results presented are for hardware (P1) without the software accelerator. Mosaic does not require custom kernels to accelerate inference. Compression is the file size compression of the LLM weights. While GPTQ sets weights to the target bit size, activations are still 16 bits during inference, which has the same memory usage as the unquantized/unpruned (dense) model. For Mosaic, the pruning target proportionally decreases inference memory usage.
Appendix G Additional Calibration Datasets and Models
Activation-based pruning and quantization methods typically use the English subset of the C4 dataset (c4-en) and the LLaMA family of LLMs for evaluation. While C4 provides broad linguistic coverage and has been shown to generalize well across a variety of downstream tasks, it may not fully capture the needs of models trained for non-English languages. For instance, a Mistral-7B-v0.1 extended with a tokenizer designed to support Japanese characters, and fine-tuned for Japanese-language benchmarks [55]. Intuitively, using a calibration dataset composed primarily of English samples may not yield a pruned model that performs optimally on Japanese reasoning tasks such as JSQuAD, JCS, JNLI, and MARC-ja [55]. To evaluate this, Table 14 compares the performance of Mistral-7B-v0.1 pruned with Mosaic using both the English (c4-en) and Japanese (c4-ja) subsets of C4. The model is evaluated on WikiText-2 (English) and four Japanese-language benchmarks covering a range of task types: reading comprehension (JSQuAD), multiple choice (JCS, JNLI), and text classification (MARC-ja). The results show that both calibration sets produce comparable perplexities and accuracies across all evaluation datasets. On average, c4-en slightly outperforms c4-ja in terms of mean accuracy and perplexity, particularly on JSQuAD and JNLI. However, c4-ja achieves marginally better performance on JCS and MARC-ja. These differences may be attributed to the significantly smaller size and narrower linguistic diversity of c4-ja (0.8 TB), compared to the larger and more heterogeneous c4-en dataset (10 TB), allowing for this dataset to perform well in multilingual settings.
@@ Appendix G Additional Calibration Datasets and Models
| Calibration Dataset | WikiText-2 | JSQuAD | JCS | JNLI | MARC-ja | Mean |
| - (Base Model) | 5.47 | 78.97 | 84.18 | 56.82 | 96.39 | 79.09 |
| c4-en | 7.21 | 72.11 | 68.01 | 29.87 | 89.12 | 64.78 |
| c4-ja | 7.26 | 70.08 | 69.52 | 27.65 | 91.50 | 64.69 |